# Measurement as Bricolage: Examining How Data Scientists Construct Target Variables for Predictive Modeling Tasks

### v1

@@ 4.2.2. Composing
P8 designed a regression model that predicted GPA given test scores, creating a model that accounted for the gap between them to capture “work ethic.” A negative residual indicated that students had lower grades than predicted by their test scores, thereby signaling work ethic concerns. While teaching an introductory freshman class, P8 was surprised to find that students predicted to have poor work ethic were more willing than others to copy homework. Her model’s predictions also closely aligned with students’ long-term persistence outcomes (e.g., retention and graduation). These observations provided evidence for the consequential validity of their model’s operationalization of “work ethic.”
@@
P2 found it easier to justify their model when it targeted a single outcome based on what software engineers and product managers viewed as “hard data.” As a result, while P2 acknowledged validity benefits with latent variable models, they did not use them in industry practice. This tension between validity and simplicity highlights how bricolage solutions often bear the mark of compromise, as data scientists negotiate between theoretical ideals and practical constraints.
@@ 4.2.4. Bridging
P6 validated their surrogate model by checking its performance against the test scores they did have available. Once they were confident that the model had “reasonable” validity, they used it to evaluate whether content changes helped students predicted to be poor readers. They found that the content updates offered a large readability improvement and deployed them on the platform, declaring the project “a really big success.” P6 acknowledged that both the gold standard tests and the formulation were imperfect. However, given the analytics focus of the project, P6 had some tolerance for validity imperfections in their bridging formulation. This tolerance for imperfection, balanced against practical gains, exemplifies the bricoleur’s pragmatic approach to problem-solving under resource constraints.
@@ 4.2.5. Refining
While constructing a target variable, participants also iteratively tweaked and refined their formulation by making small changes. This strategy involved refining: modifying the target variable definition in response to a defect along one or more criteria. In contrast to the larger change in outcomes involved with swapping, participants used refining to make more granular adjustments to their target variable.
For example, P13 used refining to narrow in on definitions for animal diseases in her dataset. While human healthcare had a mature ICD-9 disease classification scheme, P13 explained that animal health conditions remained vague and poorly specified. Clinicians were also unsure of how to map their understanding of pet diseases into a label definition: or a systematic set of classification rules in data. P13 addressed this issue by refining the label definition with clinicians.
@@
At each stage of refinement, P13 and the vet used written clinical notes about the pet to identify animals who were labeled incorrectly under the current definition. P13 then revised the label definition to correctly categorize the misclassified instances (e.g., adding a new rule-based white blood cell count). These iterations continued 20 to 30 times until the label definition sufficiently captured the disease. P13 emphasized a need for practicality during the process: they were fine with 10% of diseases being mislabeled if the major dimensions of the disease definition were covered. P13’s acceptance of imperfection while ensuring “major dimensions” were covered typifies the bricoleur’s focus on creating solutions that are “fit for purpose” rather than ideal.
@@ 4.3. Evaluating the Validity of a Problem Formulation
While applying strategies, data scientists often took a step back to evaluate the validity of their problem formulation. Drawing on domain knowledge, they engaged in theory-driven practices to probe conceptual relationships between their target variable and theoretical constructs related to their problem formulation. However, data scientists also treated their problem formulation as a material object to be scrutinized by inspecting concrete predictions generated on specific data instances. Lévi-Strauss calls this more tangible approach “the science of the concrete” — whereby the bricoleur understands their artifact through direct manipulation of material elements as opposed to abstract, conceptual reasoning. We now describe these theory-driven and data-driven practices in detail, as summarized in Table 3.
@@ 4.3.1. Theory-Driven Evaluation Practices
While engaging in theory-driven practices, data scientists used their domain knowledge to reason about a problem formulation conceptually, without scrutinizing a specific data sample. These theory-driven practices were informed by participants’ prior modeling experience and an understanding of the sociotechnical context (e.g., hospitals, classrooms) in which their models operate. Though participants had varying levels of formal measurement theory training, many theory-driven practices connect to established construct validity sub-criteria.
@@
Several participants also used multiple outcomes to develop a more complete understanding of how a construct should be conceptualized. P2 wanted to understand whether students who used hints were “gaming” the lesson – i.e., using hints to finish questions without seriously engaging with the material. Accuracy did not provide enough information to identify gaming from hint use because students could answer a question incorrectly, even after using a hint and applying effort. However, P2 found that students took longer to respond when they were engaged with the lesson. As a result, a combination of accuracy, response time, and hint use was necessary to triangulate gaming behavior. [^4] Participants who reasoned across multiple outcomes sometimes suggested composing. However, in some cases, participants used multiple outcomes to evaluate a time-lagged prediction model without explicitly leveraging multiple outcomes in the predictive model when composing wasn’t feasible (e.g., due to simplicity concerns). Reasoning across multiple outcomes connects to an assessment of content validity, or “the extent to which an operationalization wholly and fully captures the substantive nature of the construct purported to be measured” (Jacobs and Wallach, 2021).
@@
Finally, participants described their choice of outcome variable as closely linked to the intervention being informed by a predictive model. For example, P6 needed to decide which outcome time horizon (i.e., question lesson, week, semester, or year) was most appropriate for measuring the affect of platform content changes on student learning. P6 shared that “to get in the right neighborhood, where would we reasonably expect to move the needle?” If the intervention included many content changes, they leaned towards an end-of-year student performance outcome. However, if a change to the platform was more granular, they would start at a more fine grained question or lesson level. This assessment of the relationship between interventions and outcomes connects to consequential validity, or “the real-world consequences of using the measurements obtained from a measurement model” (Jacobs and Wallach, 2021).
@@ 4.3.2. Data-Driven Evaluation Practices
Several participants applied spot checks and heuristics to test whether measurements were inline with expected theoretical properties. For example, while evaluating new test items, P6 plotted students’ scores as a function of their number of question attempts. A “nice, monotonically increasing curve” with a “logarithmic flavor” provided evidence that the item encoded knowledge of a single skill. P9 checked that the base rate of a disease derived from diagnostic codes matched the expected population prevalence. A large mismatch was a signal that piggybacking off of existing disease definitions was infeasible. Participants described these heuristics as imperfect; a necessary but insufficient condition for demonstrating validity. These checks and heuristics are an example of evaluating face validity, or the degree to which an operationalization measures what it purports to measure based on a quick, surficial assessment (Coston et al., 2023).
@@ 5.1. Data Science as Bricolage
References to data science often invoke engineering metaphors: data scientists “engineer” features, “build” models, and “architect” computational solutions. Yet our findings suggest that target variable construction more closely resembles bricolage than engineering—a mode of knowledge production whereby the bricoleur makes do with the resources at hand rather than acquiring materials for a predefined plan. Our findings highlight how this process of making do unfolds in real-world projects, as data scientists re-appropriate data fields originally collected for entirely different purposes to accomplish their measurement goals. Bricolage describes how data scientists orient themselves towards measurement problems, as well as nature and quality of their solutions.
@@ 5.1.1. The Data Science Repertoire
The bricoleur operates with an inventory of odds and ends called a “repertoire.” Data scientists’ repertoire includes both material resources (e.g., datasets, computational tools) and immaterial resources (e.g., domain theories, methods) brought to bear on a modeling task. The repertoire shapes the space of possibilities that bricoleurs imagine while confronting new problems. As Louridas (1999) observes while drawing connections between bricolage and design, “the bricoleur asks his collection, whereas the engineer, like the scientist, asks the universe.” As one participant noted, a fundamental limitation is that data scientists cannot generate information that was never collected in the first place. The constrained nature of data scientists’ repertoire forces them to re-examine their limited materials from new perspectives.
In particular, materials in data scientists’ repertoire function as what Lévi-Strauss calls “signs”, or concrete objects that represent abstract concepts, but remain constrained by their imagined uses. P6’s creative re-purposing of raw user interactions in an online learning platform as a proxy for “reading ability” exemplifies the use of a sign, by unexpectedly transforming a mundane data field into an operationalization of a construct. Yet, as Levi-Strauss (1966) notes, “the possibilities remain always limited by the particular history of each piece, and by what is predetermined in it due to the original usage for which it was conceived.” Had P6 not been aware of this proxy for “reading ability”, established through a chance encounter with a colleague’s work, they may have never made this connection.
@@ 5.1.2. Engaging in Dialogue with Data
Dialogue marks a notable shift from the top-down planning approach to measurement. Under top-down planning, a practice better described as a monologue (Turkle, 2011), the data scientist defines a theoretical construct of interest, designs an operationalization, then collects data. Yet this monologue supposes that the data scientist can navigate the space of constraints and contingencies in their formulation through mental simulation, without a specific data sample or model. Our findings underscore that this is rarely the case. Instead, data scientists required tangible interactions with data to understand its affordances and constraints. It was through this very process of tangible interaction with data, borne out of resource constraints, that data scientists made creative leaps in their projects. As P6 explained, had a “magical oracle” granted their team access to all the data they required, they would have simply collected a comprehensive dataset with student test outcomes. It was the combination of resource constraints and tangible interaction with data that drove P6’s team to adaptively bridge using the limited data at hand.
@@ 5.1.3. The Emerging Artifact
Beyond constraints, Lévi-Strauss notes a personal dimension to the artifact that emerges from bricolage: “he [the bricoleur] ‘speaks’ […] through the choices he makes among the limited possibilities, the character and the life of the creator.” P8’s creative use of composing, which conceptualizes “academic work ethic” as the difference between where a student is at and where they should be, may have never emerged had they not also observed real students in real classrooms as an educator. Similarly, P12’s project, which had strong buy-in from multiple stakeholders and momentum towards completion, never came to fruition due to a validity flaw spotted by P12. Yet it was through P12’s personal interest in adjacent bodies of biostatistics literature, and their attention to methodological rigor in problem formulation, that they became aware of this otherwise invisible flaw. Each of these examples illustrates that the journey through forking paths in problem formulation depends greatly on the knowledge and experience of the bricoleur.
In sum, bricolage acknowledges that data scientists must work within the confines of their existing resources and organizational processes. These practical considerations are neither centered nor prioritized in the conventional top-down approach to measurement. Thus, short of dramatically expanding the available resources for a project, a rigid enforcement of top-down planning is unlikely to translate to a desirable project outcome. Therefore, rather than recommending more stringent enforcement of the traditional top-down measurement approach, we suggest more effectively scaffolding existing bricolage practices. We now revisit traditional top-down measurement interventions with this insight in mind, before envisioning new ways of scaffolding data science practice.
@@ 5.2. Design Implications: Reconciling Top-Down Measurement Interventions with Bricolage Practice
Pre-registration is one top-down intervention that is widely recommended in the quantitative social sciences (Van’t Veer and Giner-Sorolla, 2016). This protocol discourages data-dependent decision-making by requiring an analyst to specify outcome measures and hypotheses before running analyses. Recent work argues for the adoption of pre-registration in predictive modeling contexts (Hofman et al., 2023) — i.e., by requiring data scientists to finalize their problem formulation before training a model. Yet requiring such pre-registration is akin to blindfolding the bricoleur — it restricts the very bottom-up, iterative process that data scientists require to make do under resource constraints. Indeed, data scientists interviewed by Hofman et al. (2023) worried that pre-registration might limit their creativity, or make it challenging to later tweak their formulation if they started off with limited familiarity with their data.
@@ 5.2.2. Latent Variable Models
Researchers in the quantitative social sciences have also developed statistical tools to improve the validity of models. For example, latent variable models (e.g., (Bishop, 1998; Bartholomew et al., 2011)) use statistical machinery to characterize the relationship between unobservable theoretical constructs and observed proxies. Recent work has proposed latent variable models (e.g., (Milli et al., 2021)) and related modeling advances (e.g., (Watson-Daniels et al., 2023; Guerdan et al., 2023a)) to improve the validity of models used for decision support and recommendation systems. Yet these sophisticated approaches require advanced statistical training, multiple outcome variables, and significant implementation time. This imposes practical barriers that impede adoption. For example, P1 shared that, while they were aware of advanced multi-outcome modeling approaches, the “extra bells and whistles” provided by these approaches made them less reliable in large-scale deployment contexts.
@@ 5.2.3. Pedagogical Interventions
Our findings also highlight opportunities for pedagogical interventions to better equip practitioners to navigate target variable construction. A key insight is that data scientists often draw on prior experience, engaging in analogical reasoning by identifying parallels between their current project and those encountered previously (Jun et al., 2022a). This highlights the value of case-based learning in data science education. Given the improvisational nature of real-world data science practice, case-based approaches enable students to analyze concrete case studies and understand how experienced data scientists make trade-offs between criteria like validity, simplicity, and resource requirements. Such case-based approaches could build upon existing protocols designed to help practitioners identify potential failure modes of AI systems during the early stages of AI system design (Saxena et al., 2025; Kuo et al., 2023; Jung et al., 2025).
Reflective practice and hands-on workshops offer complementary roles in developing data scientists’ capacity to reason about measurement and validity. This mirrors how tasks like feature engineering are often described—as a “craft”—involving domain expertise, theoretical grounding, and direct engagement with data (Duboue, 2020). Workshops focused on (re)formulation strategies—such as piggybacking, composing, swapping, bridging, and refining—can expose students to the tactical, trial-and-error process of shaping a prediction task within real-world constraints. Together, these approaches offer an integrated method for teaching the complexities of bricolage-based problem formulation.

### v2

@@ 4.2.2. Composing
P8 designed a regression model that predicted GPA given test scores, creating a model that accounted for the gap between them to capture “work ethic.” A negative residual indicated that students had lower grades than predicted by their test scores, thereby signaling work ethic concerns. While teaching an introductory freshman class, P8 was surprised to find that students predicted to have poor work ethic were more willing than others to copy homework. Their model’s predictions also closely aligned with students’ long-term persistence outcomes (e.g., retention and graduation). These observations provide evidence for the consequential validity of the model.
@@
P2 found it easier to justify their model when it targeted a single outcome based on what software engineers and product managers viewed as “hard data.” As a result, while P2 acknowledged validity benefits with latent variable models, they did not use them in practice. This tension between validity and simplicity highlights how bricolage solutions often bear the mark of compromise, as data scientists negotiate between theoretical ideals and practical constraints during problem formulation.
@@ 4.2.4. Bridging
P6 validated their surrogate model by checking its performance against the test scores they did have available. Once they were confident that the model had “reasonable” validity, they used it to evaluate whether content changes helped students predicted to be poor readers. They found that the content updates offered a large readability improvement poor readers, declaring the project “a really big success.” P6 acknowledged that both the gold standard tests and the formulation were imperfect. However, given the analytics focus of the project, P6 had some tolerance for validity imperfections in their bridging formulation. This tolerance for imperfection, balanced against practical gains, exemplifies the bricoleur’s pragmatic approach to problem-solving under resource constraints.
@@ 4.2.5. Refining
During target variable construction, participants also tweaked their formulation by making small changes. This strategy involved refining: modifying the target variable definition in response to a defect along one or more criteria. In contrast to the larger change in outcomes involved with swapping, participants used refining to make more granular formulation adjustments.
For example, P13 used refining to narrow in on definitions for animal diseases in their dataset. While human healthcare had a mature ICD-9 disease classification scheme, P13 explained that animal health conditions remained vague and poorly specified. Clinicians were also unsure of how to map their understanding of pet diseases into a label definition: or a systematic set of classification rules in data. P13 addressed this issue by refining the label definition with clinicians.
@@
At each stage of refinement, P13 and the vet used written clinical notes about the pet to identify animals who were labeled incorrectly under the current definition. P13 then revised the definition to correctly categorize the misclassified instances (e.g., by adding a new rule to the definition based white blood cell count). These iterations continued 20 to 30 times until the label definition sufficiently captured the disease. P13 emphasized a need for practicality during the process: they were fine with 10% of diseases being mislabeled if the major dimensions of the disease definition were covered. P13’s acceptance of imperfection while ensuring “major dimensions” were covered typifies the bricoleur’s focus on creating solutions that are “fit for purpose” rather than ideal.
@@ 4.3. Evaluating the Validity of a Problem Formulation
While applying strategies, data scientists often took a step back to evaluate the validity of their problem formulation. Drawing on domain knowledge, they engaged in theory-driven practices to probe conceptual relationships between their target variable and theoretical constructs of interest. However, data scientists also treated their problem formulation as a material object to be scrutinized by inspecting concrete predictions generated on specific data instances. Lévi-Strauss calls this more tangible approach “the science of the concrete” — whereby the bricoleur understands their artifact through direct manipulation of material elements as opposed to abstract, conceptual reasoning. We now describe these theory-driven and data-driven practices in detail, as summarized in Table 3.
@@ 4.3.1. Theory-Driven Evaluation Practices
While engaging in theory-driven practices, data scientists used their domain knowledge to reason about a problem formulation conceptually, without scrutinizing a specific data sample. Theory-driven practices were informed by participants’ prior modeling experience and an understanding of the sociotechnical context (e.g., hospitals, classrooms) in which their models operate. Though participants had varying levels of formal measurement theory training, many theory-driven practices connect to established construct validity sub-criteria.
@@
Several participants used multiple outcomes to develop a more complete understanding of how a construct should be conceptualized. For example, P2 wanted to understand whether students who used hints were “gaming” the lesson – i.e., using hints to finish questions without engaging with the material. Accuracy did not provide enough information to identify gaming from hint use because students could answer a question incorrectly, even after using a hint and applying effort. However, P2 found that students took longer to respond when they were engaged with the lesson. As a result, a combination of accuracy, response time, and hint use was necessary to triangulate gaming behavior. Reasoning across multiple outcomes connects to an assessment of content validity, or the extent to which an operationalization captures the substantive nature of the construct (Jacobs and Wallach, 2021).
@@
Finally, participants described their choice of outcome variable as closely linked to the intervention being informed by a model. For example, P6 needed to decide which outcome time horizon (i.e., question lesson, week, semester, or year) was most appropriate for measuring the affect of platform content changes on student learning. P6 shared that “to get in the right neighborhood, where would we reasonably expect to move the needle?” If the intervention included many content changes, they leaned towards an end-of-year student performance outcome. However, if a change to the platform was more granular, they would start at a more fine grained question or lesson level. This assessment of the relationship between interventions and outcomes connects to consequential validity, or “the real-world consequences of using the measurements obtained from a measurement model” (Jacobs and Wallach, 2021).
@@ 4.3.2. Data-Driven Evaluation Practices
Several participants applied heuristics to test whether measurements were inline with expected theoretical properties. For example, while evaluating new test items, P6 plotted students’ scores as a function of their question attempt. A “nice, monotonically increasing curve” with a “logarithmic flavor” provided evidence that the item encoded knowledge of a single skill. P9 checked that the base rate of a disease derived from diagnostic codes matched the expected prevalence. A large mismatch was a signal that piggybacking off of existing disease definitions was infeasible. Participants described these heuristics as imperfect; a necessary but insufficient condition for demonstrating validity. These checks and heuristics are an example of evaluating face validity, or the degree to which an operationalization measures what it purports to measure based on a quick, surficial assessment (Coston et al., 2023).
@@ 5.1. Data Science as Bricolage
References to data science often invoke engineering metaphors: data scientists “engineer” features, “build” models, and “architect” computational solutions. Yet our findings suggest that target variable construction more closely resembles bricolage than engineering—a mode of knowledge production whereby the bricoleur makes do with the resources at hand rather than acquiring materials for a predefined top-down plan. Our findings highlight how this process of making do unfolds in real-world projects, as data scientists re-appropriate data fields originally collected for entirely different purposes to accomplish their measurement goals. Bricolage describes how data scientists orient themselves towards measurement problems, as well as nature and quality of their solutions.
@@ 5.1.1. The Data Science Repertoire
The bricoleur operates with an inventory of odds and ends called a “repertoire.” Data scientists’ repertoire includes both material resources (e.g., datasets, computational tools) and immaterial resources (e.g., domain theories, methods) brought to bear on a modeling task. The repertoire shapes the space of possibilities that bricoleurs imagine while confronting new problems. As Louridas (1999) observes while drawing connections between bricolage and design, “the bricoleur asks his collection, whereas the engineer, like the scientist, asks the universe.” P12 eloquently translated this to data science, lamenting that “ultimately, you can’t create data that don’t exist.” The constrained nature of data scientists’ repertoire forces them to re-examine their limited materials from new perspectives to identify new and innovative solutions.
Materials in data scientists’ repertoire function as what Lévi-Strauss calls “signs”, or concrete objects that represent abstract concepts, but remain constrained by their imagined uses. P6’s creative re-purposing of raw user interactions in an online learning platform as a proxy for “reading ability” exemplifies the use of a sign, by unexpectedly transforming a mundane data field into an operationalization of a construct. Yet, as Levi-Strauss (1966) notes, “the possibilities remain always limited by the particular history of each piece, and by what is predetermined in it due to the original usage for which it was conceived.” Had P6 not been aware of this proxy for “reading ability”, established through a chance encounter with a colleague’s work, they may have never made this connection.
@@ 5.1.2. Engaging in Dialogue with Data
Dialogue marks a notable shift from the top-down planning approach to measurement. Under top-down planning, a practice better described as a monologue (Turkle, 2011), the data scientist defines a theoretical construct of interest, designs an operationalization, then collects data. Yet this monologue supposes that the data scientist can navigate the space of constraints and contingencies in their formulation through mental simulation, without a specific data sample or model. Our findings underscore that this is rarely the case. Instead, data scientists required tangible interactions with data to understand its affordances. It was through this very process of tangible interaction with data, borne out of resource constraints, that data scientists made creative leaps in their projects. As P6 explained, had a “magical oracle” granted their team access to all the data they required, they would have simply collected a comprehensive dataset with test outcomes. It was the combination of resource constraints and tangible interaction with data that drove P6 to bridge using the limited data at hand.
@@ 5.1.3. The Emerging Artifact
Beyond constraints, Lévi-Strauss notes a personal dimension to the artifact that emerges from bricolage: “he [the bricoleur] ‘speaks’ […] through the choices he makes among the limited possibilities, the character and the life of the creator.” P8’s creative use of composing, which conceptualizes “academic work ethic” as the difference between where a student is at and where they should be, may have never emerged had they not also observed real students in real classrooms as an educator. Similarly, P12’s project, which had strong buy-in from multiple stakeholders and momentum towards completion, never came to fruition due to a validity flaw spotted by P12. Yet it was through P12’s personal interest in adjacent bodies of biostatistics literature, and their attention to methodological rigor in problem formulation, that they became aware of this otherwise invisible flaw. As these experiences illustrate, the final problem formulation depends greatly on the experience of the bricoleur.
In sum, bricolage acknowledges that data scientists must work within the confines of their existing resources. This practical consideration is neither centered nor prioritized in the conventional top-down approach to measurement. Thus, short of dramatically expanding the available resources for a project, a rigid enforcement of top-down planning is unlikely to translate to a desirable project outcome. Therefore, rather than recommending more stringent enforcement of the traditional top-down measurement approach, we suggest more effectively scaffolding existing bricolage practices. We now revisit traditional top-down measurement interventions with this insight in mind, before envisioning new ways of scaffolding data science practice.
@@ 5.2. Design Implications: Reconciling Top-Down Measurement Interventions with Bricolage Practice
Pre-registration is one top-down intervention that is widely recommended in the quantitative social sciences (Van’t Veer and Giner-Sorolla, 2016). This protocol discourages data-dependent decision-making by requiring an analyst to specify outcome measures and hypotheses before running analyses. Recent work argues for the adoption of pre-registration in predictive modeling contexts (Hofman et al., 2023) — i.e., by requiring data scientists to finalize their problem formulation before training a model. Yet requiring such pre-registration is akin to blindfolding the bricoleur — it restricts the very bottom-up, iterative process that data scientists require to make do under resource constraints. Indeed, data scientists interviewed by Hofman et al. (2023) worried that pre-registration might limit their creativity, or make it challenging to later tweak their formulation in response to new challenges.
@@ 5.2.2. Latent Variable Models
Researchers in the quantitative social sciences have also developed statistical tools to improve the validity of models. For example, latent variable models (e.g., (Bishop, 1998; Bartholomew et al., 2011)) use advanced statistical machinery to characterize the relationship between unobservable theoretical constructs and observed proxies. Recent work has proposed latent variable models (e.g., (Milli et al., 2021)) and related modeling advances (e.g., (Watson-Daniels et al., 2023; Guerdan et al., 2023a)) to improve the validity of models used for decision support and recommendation systems. Yet these sophisticated approaches require advanced statistical training, multiple outcome variables, and significant implementation time. This imposes practical barriers to adoption. For example, P1 shared that, while they were aware of advanced multi-outcome modeling approaches, their “extra bells and whistles’ made them unreliable in production.
@@ 5.2.3. Pedagogical Interventions
Our findings also highlight opportunities for pedagogical interventions to better equip practitioners to navigate target variable construction. A key insight from our interviews is that data scientists often draw on prior experience, engaging in analogical reasoning by identifying parallels between their current project and those encountered previously. This highlights the value of case-based learning in data science education. Case-based approaches enable students to analyze concrete case studies and understand how experienced data scientists make trade-offs between criteria like validity, simplicity, and resource requirements. Such case-based approaches could build upon existing protocols designed to help practitioners identify potential failure modes of AI systems during the early stages of AI system design (Saxena et al., 2025; Kuo et al., 2023; Jung et al., 2025).
Reflective practice and hands-on workshops offer complementary roles in developing data scientists’ capacity to reason about validity. This mirrors how tasks like feature engineering are often described—as a “craft”—involving domain expertise, theoretical grounding, and direct engagement with data (Duboue, 2020). Workshops focused on (re)formulation strategies—such as piggybacking, composing, swapping, bridging, and refining—can expose students to the tactical, trial-and-error process of shaping a prediction task within real-world constraints. Together, these approaches offer an integrated method for teaching the complexities of bricolage-based problem formulation.